one third of the sum of two consecutive odd integers is 5 less than the greater of the two. What are the two odd integers?
Answer by checkley79(3341)

LET X & X+2 BE THE 2 CONSECUTIVE ODD INTEGERS.
(X+X+2)/3=(X+2)-5
(2X+2)/3=X-3 CROSS MULTIPLY.
3(X-3)=2X+2
3X-9=2X+2
3X-2X=2+9
X=11 ANS. FOR THE SMALLER INTEGER.
11+2=13 ANS. FOR THE LARGER INTEGER.
PROOF:
(11+13)/3=13-5
24/3=8
8=8